OpenStudy (anonymous):

if f(x)=(x-1)sin x, find f' (0)

OpenStudy (anonymous):

Use product rule, where u = x-1 and v = sinx

OpenStudy (anonymous):

f(x) = uv f'(x)= (du*v)+(dv*u) f'(x) = (1)(sin x) + (x-1)(cosx) f'(x)= sin(x) + xcos(x) - cos(x)

OpenStudy (isaiah.feynman):